# UltraViCo: Breaking Extrapolation Limits in Video Diffusion Transformers

### v1

@@ C.2 More Implementation Details
The NoRepeat Score implemented in RIFLEx (zhao2025riflex) is only applicable when the content repeats once, which makes it unsuitable for longer extrapolation tasks. We therefore modify it accordingly. Specifically, the computation of the NoRepeat Score consists of two steps: static-video filtering and repeated-frame ratio calculation. In the first step, we uniformly sample 8 frames across the video; if the mean pairwise $L_{2}$ distance among them falls below a threshold, the video is considered static and discarded. This prevents completely static videos from interfering with subsequent repetition detection. In the second step, we measure the ratio of repeated frames to the total frame count, which defines the NoRepeat Score. Following RIFLEx, we first search around the dominant-frequency period for the frame with the minimal $L_{2}$ distance to the first frame. This frame is then taken as the start of a candidate repeated sequence. We then compare each frame in this candidate sequence with the corresponding frame at the beginning of the video; frames whose $L_{2}$ distance is below the threshold are counted as repetitions. Empirically, a threshold of 55 was found to align better with human perception and was consequently applied to both steps. Finally, we report the mean NoRepeat Score across all videos as the final result. The detailed implementation code is included in the supplementary material.
@@
Since Wan does not exhibit content repetition, it is not applicable to determine the dominant frequency from the repetition period as performed in zhao2025riflex. Instead, following Sec. 3.2.1, we take the largest-amplitude frequency $\phi_{0}$ as the dominant frequency.
@@ C.3 Additional Experiments of Different Extrapolation Ratios and Models
In this section, we provide some additional extrapolation ratios from $s=2$ to $5$ and models based on 25 prompts from VBench (huang2024vbench). To evaluate the generality of UltraViCo, we test $2\times$ extrapolation on HunyuanVideo, Wan, and CogVideoX, as well as $3\times$ and $4\times$ extrapolation on CogVideoX. In addition, we assess $5\times$ extrapolation on HunyuanVideo. For Wan, we set $\alpha=0.9$. For HunyuanVideo, we use $\gamma=4$ across all ratios, with $\alpha=0.95,\beta=0.6$ at $2\times$ and $\alpha=0.9,\beta=0.8$ at $5\times$. For CogVideoX, we use $\gamma=1$ and $\beta=0.6$ for all ratios, with $\alpha=0.9$ at $2\times$ and $3\times$, and $\alpha=0.85$ at $4\times$. The configurations of other baselines follow Sec. 4.1.
@@ C.5 Acceleration of UltraViCo via Sparse Attention and Distillation
Building upon recent advances in sparse-attention-based video acceleration and distillation (fastvideo2024), UltraViCo achieves about $16\times$ speed-up without compromising performance (see Table 7).
@@ C.6 Runtime and Memory Cost
As shown in Table 8, built on top of FlashAttention (dao2022flashattention) and SageAttention (zhang2024sageattention; zhang2024sageattention2), UltraViCo incurs almost no additional overhead in either latency or memory usage.
@@ D.2 Ablation on hyperparameters
Model / Method
Time (s / iter)
Memory (per GPU)

### v2

@@ C.2 More Implementation Details
The NoRepeat Score implemented in RIFLEx (Zhao et al., 2025) is only applicable when the content repeats once, which makes it unsuitable for longer extrapolation tasks. We therefore modify it accordingly. Specifically, the computation of the NoRepeat Score consists of two steps: static-video filtering and repeated-frame ratio calculation. In the first step, we uniformly sample 8 frames across the video; if the mean pairwise $L_{2}$ distance among them falls below a threshold, the video is considered static and discarded. This prevents completely static videos from interfering with subsequent repetition detection. In the second step, we measure the ratio of repeated frames to the total frame count, which defines the NoRepeat Score. Following RIFLEx, we first search around the dominant-frequency period for the frame with the minimal $L_{2}$ distance to the first frame. This frame is then taken as the start of a candidate repeated sequence. We then compare each frame in this candidate sequence with the corresponding frame at the beginning of the video; frames whose $L_{2}$ distance is below the threshold are counted as repetitions. Empirically, a threshold of 55 was found to align better with human perception and was consequently applied to both steps. Finally, we report the mean NoRepeat Score across all videos as the final result. The detailed implementation code is included in the supplementary material.
@@
Since Wan does not exhibit content repetition, it is not applicable to determine the dominant frequency from the repetition period as performed in Zhao et al. (2025). Instead, following Sec. 3.2.1, we take the largest-amplitude frequency $\phi_{0}$ as the dominant frequency.
@@ C.3 Additional Experiments of Different Extrapolation Ratios and Models
In this section, we provide some additional extrapolation ratios from $s=2$ to $5$ and models based on 25 prompts from VBench (Huang et al., 2024). To evaluate the generality of UltraViCo, we test $2\times$ extrapolation on HunyuanVideo, Wan, and CogVideoX, as well as $3\times$ and $4\times$ extrapolation on CogVideoX. In addition, we assess $5\times$ extrapolation on HunyuanVideo. For Wan, we set $\alpha=0.9$. For HunyuanVideo, we use $\gamma=4$ across all ratios, with $\alpha=0.95,\beta=0.6$ at $2\times$ and $\alpha=0.9,\beta=0.8$ at $5\times$. For CogVideoX, we use $\gamma=1$ and $\beta=0.6$ for all ratios, with $\alpha=0.9$ at $2\times$ and $3\times$, and $\alpha=0.85$ at $4\times$. The configurations of other baselines follow Sec. 4.1.
@@ C.5 Acceleration of UltraViCo via Sparse Attention and Distillation
Building upon recent advances in sparse-attention-based video acceleration and distillation (Team, 2024b), UltraViCo achieves about $16\times$ speed-up without compromising performance (see Table 7).
@@ C.6 Runtime and Memory Cost
As shown in Table 8, built on top of FlashAttention (Dao et al., 2022) and SageAttention (Zhang et al., 2024b; a), UltraViCo incurs almost no additional overhead in either latency or memory usage.
@@ D.2 Ablation on hyperparameters
Model / Method Time (s / iter) Memory (per GPU)